Comerpickles:

What is the approximate area of a circle whose diameter is 12 cm? Use 3.14 forππ . Question 1 options: 452.16 113.04 37.68 18.84

Comerpickles:

@Vocaloid

Vocaloid:

first calculate radius = (1/2) * diameter then calculate area = pi * r^2 using 3.14 as pi

Comerpickles:

Whats the diameter?

Vocaloid:

it's in the problem

Comerpickles:

Also how would I set that up? XD

Vocaloid:

you are given the diameter plug that into radius = (1/2)*diameter to find the radius then calculate the area

Comerpickles:

(1/2) * d * r^2?

Vocaloid:

radius = (1/2) * diameter = ?

Comerpickles:

So I just write diameter?

Comerpickles:

No symbol representing it?

Vocaloid:

you are given the diameter value in the problem take that number and plug that into the equation radius = (1/2) * diameter

Comerpickles:

Ohhhhh, sorry, took me a while and I had to go look back up and see what the diameter was.

Comerpickles:

6

Vocaloid:

good, then that's the radius now find area = pi * r^2 = your answer

Comerpickles:

113.04. Sorry, you're probably tired of dealing with stupid me

Vocaloid:

it's fine that's correct

Comerpickles:

I'll try on my own more